OpenStudy (anonymous):

Can someone pleaseeee help me with these! Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4) Question 2 Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9 Question 3 Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same

OpenStudy (anonymous):

the line that passes through the points (x1, y1) and (x2, y2): \[y - y1 = \left( \frac{ y2 - y1 }{ x2 - x1 } \right) \left( x-x1 \right)\]

OpenStudy (anonymous):

Q2. Take 3 on right hand side Q3. Take 4 on right hand side

OpenStudy (anonymous):

I got question 2's answer I just don't get one and 3